OpenStudy (priyar):

find the area bounded by the parabola (y-2)^2=x-1, the tangent at the point (2,3) and the x-axis .

OpenStudy (priyar):

pls integrate wrt y so that i can check it with my answer...

OpenStudy (priyar):

@ganeshie8 @freckles @UnkleRhaukus

OpenStudy (anonymous):

Find area under parabola from x =0 to x=1 Find eqaution of the tangent Find the area under tangent from its x intersept to x=1 Subtract area under tangent from area under parabola Area under parabola y = 4x^2 This parabola passes through (0,0) Definite Integrate from x =0 to point of tangent x= 1 Integral y = (4x^3)/3 Definite Integrate from x =0 to point of tangent x= 1 = 4 x 13/3 – 0 = 4/3 units Equation of tangent line Slope of parabola at (1,4) Differentiate y = 4x^2 dy/dx = 8x = slope slope at x=1 would be 8 x1 = 8 equation of tangent y = mx + b y = 8x + b since this line passes through (1,4) 4 = 8 x 1 + b b = -4 tangent equation y= 8x – 4 x intercept at y=0 0= 8x -4 X=1/2 Area under tangent from x= ½ to x=1 Integrate y = 8x - 4 Integral y = (8x^2)/2 – 4x Definite integral from x= ½ to x= 1 ((8 x 1)/2 – 4 x 1) – ((8 x ½ 2/2) – 4x1/2) 0 – (-1) = 1 unit Subtract area under tangent from area under parabola 4/3 – 1 = 1/3 units – Answer :)

OpenStudy (priyar):

this question is unrelated to my question.. @Thehulk49 ??

OpenStudy (freckles):

I like to graph my equations first: \[y=\frac{x}{2}+2 \\ (y-2)^2=x-1 \\ \\ y=0 \\ .... \\ \text{ solving for intersections } \\ (\frac{x}{2}-2)^2=x-1 \\ \frac{x^2}{4}-2x+4=x-1 \\ x^2-8x+16=4x-4 \\ x^2-12x+20=0 \\ (x-10)(x-2)=0 \\ x=10, 2 \\ \]  if I didn't make a mistake in finding the tangnet line or finding the intersections